# Efficient and Comprehensive Feature Extraction in Large Vision-Language Model for Clinical Pathology Analysis

### v2

@@ Abstract
Pathological diagnosis is vital for determining disease characteristics, guiding treatment, and assessing prognosis, relying heavily on detailed, multi-scale analysis of high-resolution whole slide images (WSI). However, traditional pure vision models face challenges of redundant feature extraction, whereas existing large vision-language models (LVLMs) are limited by input resolution constraints, hindering their efficiency and accuracy. To overcome these issues, we propose two innovative strategies: the mixed task-guided feature enhancement, which directs feature extraction toward lesion-related details across scales, and the prompt-guided detail feature completion, which integrates coarse- and fine-grained features from WSI based on specific prompts without compromising inference speed. Leveraging a comprehensive dataset of 490,000 samples from diverse pathology tasks—including cancer detection, grading, vascular and neural invasion identification, and so on—we trained the pathology-specialized LVLM, OmniPT. Extensive experiments demonstrate that this model significantly outperforms existing methods in diagnostic accuracy and efficiency, offering an interactive, clinically aligned approach for auxiliary diagnosis in a wide range of pathology applications.
Figure 1: From top to bottom are examples of our OmniPT, a vision-language model optimized for pathology, applied to referring expression detection, segmentation, and visual question answering. Notably, in the first example, OmniPT is tasked with detecting cancer cell nuclei within blood vessels. Results show that OmniPT accurately identifies most nuclei within vessels without mistakenly detecting any outside, demonstrating its capability to understand pathological concepts and reason effectively.
@@ 1 Introduction
In the field of digital pathology, artificial intelligence is playing an increasingly vital role [16, 15, 52, 54]. Many pure vision deep learning models have been developed to assist in the diagnosis of pathology WSIs [10, 11, 51, 45, 46]. With the rapid development of large language models (LLM) [5, 34, 1, 43, 44, 14] and large vision-language models (LVLM) [30, 4, 8, 26, 12], these models have shown substantial auxiliary capabilities across various domains, and recent research [40, 41, 36, 31, 2, 13, 39, 37] has attempted to apply LVLM to pathology. However, existing methods still face significant limitations: traditional pure vision models [10, 11, 51, 45, 46] require WSIs to be divided into thousands of patches, with an encoder network extracting features from each patch and an aggregator network synthesizing the final result. These methods inevitably extract numerous redundant features, leading to a prolonged diagnostic process. Current pathological LVLMs [40, 41, 36, 31, 2, 13, 39, 37, 49], due to input constraints, can only process either single pathological image patches or low-resolution thumbnails of WSI. While this improves processing speed, it either lacks global information for single patches or loses substantial detail information for WSI, making it difficult to meet clinical assisted diagnostic requirements. Furthermore, experimental analysis (given in Section 3) of image tokens that LVLM focuses on for decision-making reveals that existing LVLMs often overly emphasize the features of a few key tokens in the input image. While this feature extraction pattern can summarize image content, it fails to comprehensively capture multi-scale features related to lesions, thus affecting the model’s diagnostic performance.
@@
To enhance the accuracy and reliability of intelligent pathological diagnosis and analysis, we have developed an efficient and comprehensive feature extraction scheme specifically tailored for LVLM in pathology, providing complete, multi-scale feature support for various types of pathological diagnostic analysis tasks. To address the issue of the model focusing only on a few key tokens of the image, we introduce the mixed task-guided feature enhancement (MTGFE) strategy. Through adding instruction-following data for detecting and segmenting diverse pathological concepts, coupled with corresponding model module improvements, we enhance the model’s ability to perceive lesion-related detailed features across the whole image while achieving full coverage of visual task types in pathological analysis. Furthermore, given the need for multi-scale features in pathology slide analysis, we designed the prompt-guided detail feature completion (PGDFC) strategy. This strategy first captures the coarse-grained global features of a WSI and then, based on specific task requirements provided by prompts, extracts fine-grained features from key focus areas. By merging coarse- and fine-grained features, this approach enhances accuracy across tasks while avoiding the input of exhaustive detail features, thereby maintaining high inference speed.
@@
To make LVLM truly applicable to clinical pathology for auxiliary diagnosis, we gathered visual instruction-following data for multiple organs and tasks from several institutions, based on diagnostic items in actual pathology reports. This data trains the model to meet the diverse clinical needs of pathologists through human-computer interaction. These tasks include cancer region detection and segmentation, cancer grading and classification, identification of vascular and neural invasion, and lymph node metastasis detection, among others, with each task containing meticulously compiled datasets ranging from hundreds to thousands of instruction-following data samples. Furthermore, to strengthen the model’s understanding of foundational pathology concepts, we collected training data for fundamental tasks such as nucleus detection and classification, vascular and neural detection, lymph node detection, and tumor-infiltrating lymphocyte identification. Additionally, we integrated pathology image-text data from publicly available online resources, including the PubMed database [17, 40], pathology textbooks and atlases [20, 40], The Cancer Genome Atlas (TCGA) [39, 9], Twitter posts [21, 41], and educational histopathology videos on YouTube [22, 36]. This integration yielded a comprehensive dataset covering 20 organs with approximately 490,000 training samples. Leveraging our efficient and comprehensive feature extraction scheme and this extensive dataset, we trained OmniPT, a specialized LVLM for pathology, capable of providing comprehensive pathology auxiliary diagnostic services through human-computer interaction. Extensive experimental results demonstrate that OmniPT outperforms existing pathology LVLMs across a range of diagnostic tasks, better aligning with the actual needs of clinical practice.
@@
  • We propose the mixed task-guided feature enhancement strategy, which guides large vision-language model to capture detailed features in pathology images by constructing fine-grained tasks focused on local image features and integrating corresponding improvements in model architecture. This effectively addresses the model’s tendency to rely excessively on global features represented by a few key tokens of the image.
@@
  • We integrated multi-source data to construct a pathology visual instruction tuning dataset covering 20 organs with approximately 490,000 training samples. Based on this dataset, we trained the OmniPT pathology large vision-language model. Extensive experimental validation shows that OmniPT surpasses existing pathology large vision-language models across various diagnostic tasks, aligning more closely with clinical pathology diagnostic needs through an interactive paradigm.
@@ 2 Related Work
have been applied to pathology image analysis for quite some time. Early research primarily focused on designing specialized model architectures for specific tasks, including nuclei segmentation [15], vessel segmentation [16], microvascular invasion detection [54], and cancer grading [52]. In contrast, research on overall pathology WSI analysis tasks, such as cancer subtyping, metastasis detection, and prognosis prediction, commonly adopts a multiple instance learning (MIL) approach [23, 38, 50, 42]. This approach treats an entire WSI as a “bag”, partitioning the WSI into tens of thousands of patches similar in size to conventional images, with each patch considered an “instance.” In practice, researchers typically employ an encoder network to extract features from each patch and then aggregate these features to produce an analysis result for the entire WSI.
To enhance the transferability and generalization of these methods, recent efforts have focused on self-supervised training of the patch encoder using vast amounts of unlabeled WSIs [10, 11, 51, 45, 46]. These self-supervised approaches significantly improve performance across various general tasks and bolster the model’s ability to identify rare diseases. However, this approach often requires substantial time to extract features from each patch, and critical disease-related features may be obscured by an overwhelming number of unrelated features.
@@
have also been explored for pathology image analysis in the past year [40, 41, 36, 31, 2, 13, 39, 37, 49]. These models typically utilize LLaVA-based architectures [30], fine-tuned with instruction-following data curated through various methods and sources. PathAsst [40] gathered image-caption pairs from the PubMed database and internal materials to train a CLIP [35] model, replacing LLaVA’s vision encoder, and leveraged ChatGPT [34] to transform simple image-caption data into more complex instruction-following data, which also included instructions for invoking specialized models as needed. Quilt-LLaVA [36] utilized the mouse pointer trails from speakers in YouTube tutorial videos to pinpoint the pathology concept being explained, then constructed instruction-following data using GPT-4 [1]. PathMMU [41] integrated data from multiple sources to build a pathology visual question-answering training and evaluation dataset, with seven pathology experts reviewing the test set to ensure authoritative evaluation. PathAlign [2] employed a Q-Former structure from BLIP-2 [28] to extract features from WSIs as inputs for the LLM, while PA-LLaVA [13] developed a scale-invariant connector to prevent information loss from image resizing.
However, due to structural limitations, most of these models can only accept standard-sized images, handling only single pathology image patches or low-resolution WSI thumbnails. As a result, they lack global context for single patches and lose significant detail for WSIs, posing challenges for clinical diagnostic support. Although PathAlign [2] enables full WSI input, it still requires feature extraction from each patch. Furthermore, these models are typically limited to image description tasks or visual question-answering for image classification, lacking the capability for fine-grained tasks like detection and segmentation, as well as complex reasoning tasks that require multiple diagnostic steps.
@@
3 Analysis of Drawback in Existing LVLM
3.1 Preliminaries
Today’s most prominent open-source LVLM, like LLaVA [30], successfully integrate vision and language capabilities. For input pair $x=(x_{v},x_{t})$, where $x_{v}$ represents the image and $x_{t}$ represents the text prompt, the model first processes them through two embedding networks: vision encoder $\mathcal{V}$, consisting of a CLIP-based Vision Transformer (ViT) [35] followed by a projection layer, maps the image into feature embedding $e_{v}=(e_{v}^{1},...,e_{v}^{N})$ where $N$ is the number of image tiles, while text encoder $\mathcal{T}$, comprising a tokenizer and an embedding layer, transforms the text into feature embedding $e_{t}=(e_{t}^{1},...,e_{t}^{M})$ where $M$ is the number of text tokens. Both embeddings lie in the same feature space and serve as input to the large language model $\mathcal{M}$, which generates the output sequence $y$. The model can be formalized as: $y=\mathcal{M}(e_{v},e_{t})$, where $e_{v}=\mathcal{V}(x_{v})$ and $e_{t}=\mathcal{T}(x_{t})$.
Figure 3: Overview of the proposed OmniPT. Left: the architecture of OmniPT with feature enhancement and completion strategy. Right: the detailed structure of the multi-scale feature fusion vision encoder.
@@ 3.2 Decision-Dependent Image Tokens Analysis
To further investigate the image feature patterns that LVLM relies on during the answer generation and decision-making processes, and to optimize the model to focus more effectively on task-relevant features, thereby improving accuracy in responding to human queries, we conducted an attention pattern analysis on existing medical LVLMs. Specifically, we selected two representative models, LLaVA-Med [27] and Quilt-LLaVA [36], using a unified prompt, “What cancer subtype is shown in this image?” to guide the models in performing cancer subtype identification on pathology slides. Our proposed OmniPT model was included as a comparison. Visualization of the relevant analysis results is shown in Figure 2.
We extracted the attention matrix from the input layer of $\mathcal{M}$ and averaged the attention values across all heads to obtain matrix $\Psi\in\mathbb{R}^{(N+M)\times(N+M)}$. In this matrix, the $i$-th row of $\Psi$ represents the attention distribution of the $i$-th embedding token in the input of $\mathcal{M}$ towards other tokens. Since a decoder-only Transformer model is used, $\Psi$ takes the form of a lower triangular matrix. To analyze the relationship between the upcoming generated content and the image tokens, we selected the attention values of the final embedding token towards all image tokens, denoted as $\Psi_{(N+M),e_{v}}\in\mathbb{R}^{N}$, and restored it to a two-dimensional representation of the original image. We then generated a heatmap and overlaid it on the input image for visualization, as shown in the first column of Figure 2.
In the heatmap, attention intensity is mapped from blue (low) to red (high). Cancerous regions in the input pathology slide $x_{v}$ are annotated by pathologists with green contours. Ideally, to achieve accurate cancer subtype identification, the model should focus on image features within the cancerous region rather than on other tissue and background areas. However, the heatmaps for LLaVA-Med [27] and Quilt-LLaVA [36] reveal that only a few key image tokens receive high attention weights, and these key tokens are primarily located outside the cancerous regions.
To further analyze the information encoded by these key image tokens, we selected one key image token $e_{v}^{k}$ and one ordinary token $e_{v}^{o}$ from each experiment, and visualized their attention distributions over all image tokens, denoted as $\Psi_{e_{v}^{k},e_{v}}\in\mathbb{R}^{N}$ and $\Psi_{e_{v}^{o},e_{v}}\in\mathbb{R}^{N}$, respectively. The selected key and ordinary tokens are indicated in the first column of Figure 2 with red and yellow boxes, respectively. The corresponding attention heatmap visualizations are presented in the second and third columns of Figure 2.
@@
In contrast, in the optimized OmniPT model, the heatmap of $\Psi_{(N+M),e_{v}}$ shows that key image tokens are concentrated in the cancerous regions, indicating that these areas contribute more information to $\mathcal{M}$, aiding in accurate cancer subtype identification. From the heatmap of $\Psi_{e_{v}^{k},e_{v}}$, it can be seen that although the key token maintains high attention values across all preceding image tokens, its focus on tokens within the cancerous regions is significantly higher. This demonstrates that OmniPT can more precisely capture critical diagnostic features in pathology WSIs, thus achieving a better response to user instructions. The next section will elaborate on the optimization strategies employed in OmniPT.
@@ 4 Method
To address the identified limitations of existing LVLMs, specifically their tendency to over-rely on key tokens and inability to comprehensively capture multi-scale pathological features, we propose a novel framework that enhances both feature extraction precision and efficiency. Our approach consists of two complementary strategies: the mixed task-guided feature enhancement (MTGFE) and the prompt-guided detail feature completion (PGDFC). These strategies work in concert to improve the model’s capability in pathological image analysis by targeting the core challenges revealed in our previous analysis while maintaining computational efficiency and diagnostic accuracy. OmniPT trained with these two strategies is shown in Figure 3. Below, we elaborate on these strategies and their implementation.

### v3

# Efficient and Comprehensive Feature Extraction in Large Vision-Language Model for Pathology Analysis

@@ Abstract
Pathological diagnosis is vital for determining disease characteristics, guiding treatment, and assessing prognosis, relying heavily on detailed, multi-scale analysis of high-resolution whole slide images (WSI). However, existing large vision-language models (LVLMs) are limited by input resolution constraints, hindering their efficiency and accuracy in pathology image analysis. To overcome these issues, we propose two innovative strategies: the mixed task-guided feature enhancement, which directs feature extraction toward lesion-related details across scales, and the prompt-guided detail feature completion, which integrates coarse- and fine-grained features from WSI based on specific prompts without compromising inference speed. Leveraging a comprehensive dataset of 490K samples from diverse pathology tasks, we trained the pathology-specialized LVLM, OmniPath. Extensive experiments demonstrate that this model significantly outperforms existing methods in diagnostic accuracy and efficiency, providing an interactive, clinically aligned approach for auxiliary diagnosis in a wide range of pathology applications.
@@ 1 Introduction
Figure 1: Dialogue examples of our OmniPath, a vision-language model optimized for pathology, applied to referring expression detection, segmentation, and visual question answering. Notably, in the first example, OmniPath is tasked with detecting cancer cell nuclei within blood vessels. Results show that OmniPath accurately identifies most nuclei within vessels without mistakenly detecting any outside, demonstrating its capability to understand pathological concepts and reason effectively.
@@
In the field of digital pathology, artificial intelligence is playing an increasingly vital role [feng2021edge, feng2021mutual, Zhang_2023_CVPR]. Many pure vision deep learning models have been developed to assist in the diagnosis of pathology WSIs [chen2022scaling, UNI, xu2024whole, vorontsov2024foundation, wang2024pathology]. With the rapid development of large language models (LLM) [gpt4, llama3, qwen25, deepseekr1] and large vision-language models (LVLM) [llava15, qwen2vl, shikra, lisa], these models have shown substantial auxiliary capabilities across various domains, and recent research [pathasst, pathmmu, quiltllava, pathchat, pathalign, pallava, pathgen, prism] has attempted to apply LVLM to pathology.
@@
However, existing methods still face significant limitations: traditional pure vision models require WSIs to be divided into thousands of patches, with an encoder network extracting features from each patch and an aggregator network synthesizing the final result. These methods inevitably extract numerous redundant features, leading to a prolonged diagnostic process. Current pathological LVLMs, due to input constraints, can only process either single pathological image patches or low-resolution thumbnails of WSI. While this improves processing speed, it either lacks global information for single patches or loses substantial detail information for WSI, making it difficult to meet clinical assisted diagnostic requirements. Furthermore, experimental analysis (given in Sec. 3) of image tokens that LVLM focuses on for decision-making reveals that existing LVLMs often overly emphasize the features of a few key tokens in the input image. While this feature extraction pattern can summarize image content, it fails to comprehensively capture multi-scale features related to lesions, thus affecting the model’s diagnostic performance.
@@
To enhance the accuracy and reliability of intelligent pathological diagnosis and analysis, we develop an efficient and comprehensive feature extraction scheme specifically tailored for LVLM in pathology, providing complete, multi-scale feature support for various types of pathological diagnostic analysis tasks. In addressing the issue of the model focusing only on a few key tokens of the image, we introduce the mixed task-guided feature enhancement (MGFE) strategy. Through adding instruction-following data for detecting and segmenting diverse pathological concepts, coupled with corresponding model module improvements, we enhance the model’s ability to perceive lesion-related detailed features across the whole image while achieving full coverage of visual task types in pathological analysis. Furthermore, given the need for multi-scale features in pathology slide analysis, we design the prompt-guided detail feature completion (PGFC) strategy. This strategy first captures the coarse-grained global features of a WSI and then, based on specific task requirements provided by prompts, extracts fine-grained features from key focus areas. By merging coarse- and fine-grained features, this approach enhances accuracy across tasks while avoiding the input of exhaustive detail features, thereby maintaining high inference speed.
@@
To make LVLM truly applicable to clinical pathology for auxiliary diagnosis, we curate visual instruction-following data for multiple tasks from several institutions, based on diagnostic items in actual pathology reports. These tasks include cancer region detection and segmentation, cancer grading and subtyping, identification of vascular and neural invasion, and lymph node metastasis detection, among others. Furthermore, to strengthen the model’s understanding of foundational pathology concepts, we collected training data for fundamental tasks such as nucleus detection and classification, vascular and neural detection, lymph node detection, and tumor-infiltrating lymphocyte identification. Additionally, we integrated pathology image-text data from publicly available online resources, including the PubMed database [arch, pathasst], pathology textbooks and atlases [pathvqa, pathasst], The Cancer Genome Atlas (TCGA) [pathgen, wsivqa], Twitter posts [openpath, pathmmu], and educational histopathology videos on YouTube [quilt1m, quiltllava]. This integration yielded a comprehensive dataset covering 21 organs with approximately 490K training samples. Leveraging our efficient and comprehensive feature extraction scheme and this extensive dataset, we trained OmniPath, a specialized LVLM for pathology, capable of providing comprehensive pathology auxiliary diagnostic services through human-computer interaction. Extensive experimental results demonstrate that OmniPath outperforms existing pathology LVLMs across a range of diagnostic tasks, better aligning with the actual needs of clinical practice.
@@
  • The mixed task-guided feature enhancement strategy is devised to direct large vision-language models to capture detailed pathology image features through fine-grained tasks targeting local features, while incorporating model module improvements. This approach reduces the model’s overreliance on global features represented by a few key image tokens.
@@
  • We developed OmniPath, a pathology-specific large vision-language model trained on a multi-source dataset encompassing 21 organs with 490K samples. Extensive experiments demonstrate OmniPath’s superior performance over existing models across diagnostic tasks, aligning with clinical needs through an interactive framework.
@@ 2 Related Work
have long been used in pathology image analysis, initially focusing on specialized architectures for tasks like nuclei segmentation [feng2021mutual], vessel segmentation [feng2021edge], and microvascular invasion detection [Zhang_2023_CVPR]. Broader WSI analysis tasks, such as cancer subtyping and prognosis prediction, often employ multiple instance learning (MIL) [abmil, transmil2021shao, camel2019xu, thandiackal2022differentiable], which partitions WSIs into numerous patches for feature extraction and aggregation. To enhance generalization, recent work leverages self-supervised training on large-scale unlabeled WSIs [chen2022scaling, UNI, xu2024whole, vorontsov2024foundation, wang2024pathology], improving overall performance and rare disease identification. However, this approach remains computationally intensive and risks diluting critical pathological features with irrelevant information.
@@
have recently been explored for pathology image analysis [pathasst, pathmmu, quiltllava, pathchat, pathalign, pallava, pathgen, prism, pathinsight], typically using LLaVA-based architectures [llava] fine-tuned with instruction-following data from diverse sources. PathAsst [pathasst] trained a CLIP [clip] model with PubMed and internal image-caption pairs, using ChatGPT [instructgpt] to generate more complex instructions. Quilt-LLaVA [quiltllava] extracted pathology concepts from YouTube tutorials via mouse pointer trails and constructed training data with GPT-4 [gpt4]. PathMMU [pathmmu] compiled multi-source data for pathology visual QA, with expert-reviewed test sets. PathAlign [pathalign] used a BLIP-2 [blip] Q-Former to extract WSI features for LLMs, while PA-LLaVA [pallava] introduced a scale-invariant connector to mitigate image resizing losses.
However, most models are limited to standard-sized images, processing only single patches or low-resolution WSI thumbnails, leading to a loss of global context or fine details. While PathAlign [pathalign] supports full WSI input, it still requires per-patch feature extraction. Additionally, these models are mainly suited for image description and visual QA but lack capabilities for fine-grained tasks like detection and segmentation, as well as complex multi-step diagnostic reasoning.
3 Analysis of Drawback in Existing LVLMs
3.1 Preliminaries
Today’s most prominent open-source LVLM, like LLaVA [llava], successfully integrate vision and language capabilities. For input pair $x=(x_{v},x_{t})$, where $x_{v}$ represents the image and $x_{t}$ represents the text prompt, the model first processes them through two embedding networks: vision encoder $\mathcal{V}$, consisting of a CLIP-based Vision Transformer (ViT) [clip] followed by a projection layer, maps the image into feature embedding $e_{v}=(e_{v}^{1},...,e_{v}^{N})$ where $N$ is the number of image tiles, while text encoder $\mathcal{T}$, comprising a tokenizer and an embedding layer, transforms the text into feature embedding $e_{t}=(e_{t}^{1},...,e_{t}^{M})$ where $M$ is the number of text tokens. Both embeddings lie in the same feature space and serve as input to the large language model $\mathcal{M}$, which generates the output sequence $y$. The model can be formalized as: $y=\mathcal{M}(e_{v},e_{t})$, where $e_{v}=\mathcal{V}(x_{v})$ and $e_{t}=\mathcal{T}(x_{t})$.
@@ 3.2 Decision-Dependent Image Tokens Analysis
To further investigate the image feature patterns that LVLM relies on during the answer generation and decision-making processes, and to optimize the model to focus more effectively on task-relevant features, thereby improving accuracy in responding to human queries, we conducted an attention pattern analysis on existing medical LVLMs. Specifically, we selected two representative models, LLaVA-Med [llavamed] and Quilt-LLaVA [quiltllava], using a unified prompt, “What cancer subtype is shown in this image?” to guide the models in performing cancer subtype identification on pathology slides. Our proposed OmniPath model was included as a comparison. Visualization of the relevant analysis results is shown in Fig. 2.
We extracted the attention matrix from the input layer of $\mathcal{M}$ and averaged the attention values across all heads to obtain matrix $\Psi\in\mathbb{R}^{(N+M)\times(N+M)}$. In this matrix, the $i$-th row of $\Psi$ represents the attention distribution of the $i$-th embedding token in the input of $\mathcal{M}$ towards other tokens. Since a decoder-only Transformer model is used, $\Psi$ takes the form of a lower triangular matrix. To analyze the relationship between the upcoming generated content and the image tokens, we selected the attention values of the final embedding token towards all image tokens, denoted as
$\Psi_{(N+M),e_{v}}\in\mathbb{R}^{N},$
and restored it to a two-dimensional representation with the same shape as the original image feature. We then generated a heatmap and overlaid it on the input image for visualization, as shown in the first column of Fig. 2.
In the heatmap, attention intensity is mapped from blue (low) to red (high). Cancerous regions in the input pathology slide $x_{v}$ are annotated by pathologists with green contours. Ideally, to achieve accurate cancer subtype identification, the model should focus on image features within the cancerous region rather than on other tissue and background areas. However, the heatmaps for LLaVA-Med [llavamed] and Quilt-LLaVA [quiltllava] reveal that only a few key image tokens receive high attention weights, and these key tokens are primarily located outside the cancerous regions.
To further analyze the differences in the image information encoded by these key image tokens and other ordinary image tokens, we selected one key image token $e_{v}^{k}$ and one ordinary token $e_{v}^{o}$ from each experiment, and visualized their attention distributions over all image tokens, denoted as $\Psi_{e_{v}^{k},e_{v}}\in\mathbb{R}^{N}$ and $\Psi_{e_{v}^{o},e_{v}}\in\mathbb{R}^{N}$, respectively. The selected key and ordinary tokens are indicated in the first column of Fig. 2 with red and yellow boxes, respectively. The corresponding attention heatmap visualizations are presented in the second and third columns of Fig. 2.
@@
In contrast, in the optimized OmniPath model, the heatmap of $\Psi_{(N+M),e_{v}}$ shows that key image tokens are concentrated in the cancerous regions, indicating that these areas contribute more information to $\mathcal{M}$, aiding in accurate cancer subtype identification. From the heatmap of $\Psi_{e_{v}^{k},e_{v}}$, it can be seen that although the key token maintains high attention values across all preceding image tokens, its focus on tokens within the cancerous regions is significantly higher. This demonstrates that OmniPath can more precisely capture critical diagnostic features in pathology WSIs, thus achieving a better response to user instructions. The next section will elaborate on the optimization strategies employed in OmniPath.
@@ 4 Method
To address the identified limitations of existing LVLMs, specifically their tendency to over-rely on key tokens and inability to comprehensively capture multi-scale pathological features, we propose a novel framework that enhances both feature extraction precision and efficiency. Our approach consists of two complementary strategies: the mixed task-guided feature enhancement (MGFE) and the prompt-guided detail feature completion (PGFC). These strategies work in concert to improve the model’s capability in pathological image analysis by targeting the core challenges revealed in our previous analysis while maintaining computational efficiency and diagnostic accuracy. OmniPath trained with these two strategies is shown in Fig. 3. Below, we elaborate on these strategies and their implementation.